OpenStudy (anonymous):

Consider the function f(x,y)=5x−9y+4xy−7(x^2)+6(y^2) defined in the unit square 0

OpenStudy (amistre64):

it appears to be hyperbolic ......

OpenStudy (amistre64):

we can get a vector derivative if we take the gradient: but I aint sure how that would help yet

OpenStudy (amistre64):

f(x,y)=5x−9y+4xy−7(x^2)+6(y^2) fx(x,y) = 5 +4y -14x fy(x,y) = -9 +4x +12y

OpenStudy (amistre64):

gf = (5 +4y -14x , -9 +4x +12y) can also define the normal to the surface; when this is pointing straight up or down would give you a tangent plane for a min or max right?

OpenStudy (amistre64):

straight up is the same as (0,y); so id say make x=0 for starters ... just a hunch

OpenStudy (amistre64):

gf = (5 +4y , -9 +12y)

OpenStudy (anonymous):

nice multivariable

OpenStudy (amistre64):

that mighta been a wrong step lol

OpenStudy (amistre64):

aint it tho :)

OpenStudy (amistre64):

gf = (5 +4y -14x , -9 +4x +12y) maybe when 5 +4y -14x = 0

OpenStudy (anonymous):

I don't remember how to do this.. but we dealing with a gradient mic in one of my classes

OpenStudy (anonymous):

mic check one two one two

OpenStudy (amistre64):

5 +4y -14x = 0 y = 14x-5 x = 4y +5 ----- ------ 4 14

OpenStudy (anonymous):

can you use LaTeX here?

OpenStudy (amistre64):

you can \(\try\)

OpenStudy (amistre64):

when: 5 +4y -14x = -9 +4x +12y we might get a zero derivative.... 5 + 4y -14x 9 -12y -4x ----------- 14 -8y -18x = 0 18x +8y = 14 ... i wonder...

OpenStudy (amistre64):

i was sooo close lol add the gradient parts and do something....

OpenStudy (amistre64):

i was on the right track; when the x and y parts of the gradient =0 we got a tangent plane that is flat.... to indicate a high and low

OpenStudy (amistre64):

18x +8y = 14; we can use values between 0 and 1 to find out what works now

OpenStudy (amistre64):

the line y=-(18x+14)/8 should give us the answers right? or y = -(9x+7)/4